### v1

# G2D: Boosting Multimodal Learning with Gradient-Guided Distillation

Abstract
Multimodal learning aims to leverage information from diverse data modalities to achieve more comprehensive performance. However, conventional multimodal models often suffer from modality imbalance, where one or a few modalities dominate model optimization, leading to suboptimal feature representation and underutilization of weak modalities. To address this challenge, we introduce Gradient-Guided Distillation (G2D), a knowledge distillation framework that optimizes the multimodal model with a custom-built loss function that fuses both unimodal and multimodal objectives. G2D further incorporates a dynamic sequential modality prioritization (SMP) technique in the learning process to ensure each modality leads the learning process, avoiding the pitfall of stronger modalities overshadowing weaker ones. We validate G2D on multiple real-world datasets and show that G2D amplifies the significance of weak modalities while training and outperforms state-of-the-art methods in classification and regression tasks. Our code is available here.
1 Introduction
(a) Audio Performance
(b) Video Performance
(c) Multimodal Performance
Figure 1: Performance of unimodal-only, unimodal in multimodal training, and purely multimodal models on the CREMA-D test set for multimodal classification. (a) audio modality is indifferent to training configurations; (b) video modality is vulnerable to the audio modality in a multimodal setting; (c) performance of the multimodal model is not optimal because of modality imbalance. G2D limits the optimization of superior modality and enhances the video modality to optimize the multimodal performance.
Multimodal learning is one of the most prominent multidisciplinary research areas due to the increasing demand to develop intelligent agents that perceive information from diverse sensory modalities. One of the primary challenges of multimodal learning models is the modality imbalance phenomenon [24, 21, 37, 42], also known as the modality competition [21, 17, 16] or modality laziness [46, 7]. Modality imbalance occurs when one modality dominates and other modalities are underutilized in the optimization of multimodal learning models. This causes (i) inferior multimodal performance compared to unimodal models [30, 24, 36], or (ii) a larger gap in individual modality when they are optimized jointly but still improve the model performance [21, 8]. This imbalance occurs due to poor alignment of the modalities, model overfitting to the modalities [36], and differences in the rate of model convergence [48]. An example of a modality imbalance with the multimodal dataset CREMA-D [3] is given in  Figure 1. It is evident that the audio features(Figure 1(a)) dominate the video features(Figure 1(b)) when optimized in a multimodal fashion. However, video features give better performance with unimodal training. This results in sub-par performance with the joint multimodal training, shown in  Figure 1(c). In this work, our goal is to (i) increase the performance of weak modalities in multimodal settings and (ii) increase the overall performance of the multimodal model in downstream supervised tasks.
In recent years, many methods have been proposed to address modality imbalance in multimodal learning [36, 24, 21, 39, 9, 43, 7, 8, 37, 46, 42, 16]. Gradient modulation is one of the popular approaches in state-of-the-art methods to dynamically modify multimodal optimization gradients and maximize equal contributions from all modalities. A common form of gradient modulation is dynamically increasing the gradients of weak modalities only for late fusion [24, 45, 43] or for any type of fusion methods [21, 8] during the training process. Multiple variations of these gradient modulations also exist, for example, alternating gradients of each modality [46] and controlling the dominant modality gradients [9]. Although there exist multiple gradient modulation methods, there are very few methods that optimize both unimodal and multimodal learning objectives [37, 7] to add the benefits of both worlds. We aim to follow this trend in this work and introduce a novel optimization strategy by incorporating knowledge distillation.
In this work, we aim to utilize the full potential of both unimodal and multimodal learning for the given multimodal downstream task. We propose a framework Gradient-Guided Distillation (G2D) that transfers knowledge from multiple unimodal teachers to multimodal student models. Novelty of G2D is its use of knowledge distillation with a new learning objective and gradient modulation technique to mitigate modality imbalance and produce state-of-the-art results in multimodal learning. As depicted in Figure 1, G2D improves the feature quality of multimodal encoder, allowing both the audio and video encoders to approach the accuracy of their unimodal counterparts when integrated into the multimodal model. This leads to an overall more balanced and better-performing multimodal model while addressing the modality imbalance issue (Figure 1(c)). Our three-fold contributions in this work are:
  • We introduce a knowledge distillation framework called G2D that adapts a new optimization technique to fuse both unimodal and multimodal learning to enhance the performance of downstream tasks.
  • We propose a new Sequential Modality Prioritization strategy that dynamically balances the optimization of weak and dominant modalities to mitigate imbalance.
  • With extensive experiments on six real-world datasets, we show that G2D minimizes modality imbalance and achieves superior performance. We also show that our approach is adaptive to existing methods.
2 Related Work
[…]
Knowledge distillation (KD) transfers knowledge from a larger and more complex model (teacher) to a smaller and more efficient model (student[15, 2, 31]. Multimodal KD extends this concept by leveraging information from multiple modalities to enhance learning, which supports several real-world AI applications ([40, 22, 27, 35, 44]). Multimodal KD distills knowledge from one modality to improve performance of other modalities or in multimodal downstream tasks. Multimodal KD has demonstrated substantial benefits, including improved performance, better multimodal alignment, and enhanced generalization across modalities ([18, 41, 10, 44]). Multimodal KD has been used in multiple real-world problems like medical imaging [35] to address missing modality and action recognition [27] to transfer knowledge from multimodal ensemble to a unimodal model. In this paper, we build upon recent advances in multimodal KD [7], specifically targeting the modality imbalance problem during multimodal training. By utilizing unimodal teacher models to guide the multimodal student model, we ensure balanced learning across modalities.
Several methods have been proposed to mitigate the modality imbalance through gradient modulation, feature rebalancing, or modality-specific learning rate adjustments ([24, 21, 9, 43, 45, 8]). Gradient modulation techniques dynamically adjust gradients to balance modality contributions during training [24, 21]. However, these methods often require careful tuning of hyperparameters, which can limit their generalizability. Feature rebalancing aims to optimize multimodal interaction by adjusting the contribution of each modality by enhancing the performance of unimodal learners [46, 37, 16, 38]. There is another perspective of alleviating modality imbalance with a proper label fitting using contrastive learning [42]. In the context of multimodal KD, limited work has used unimodal teachers to supervise a multimodal student. UMT [6] directly supervises with unimodal teachers, while UME [7] aggregates their logits. Choosing between them requires empirical tuning, limiting adaptability across tasks and datasets. In this work, we propose the G2D framework to address these limitations by combining multimodal knowledge distillation with a new gradient modulation technique. Unlike previous approaches that require careful hyperparameter tuning ([24, 21, 7, 6]), our approach dynamically suppresses dominant modalities based on insights from unimodal teachers. This not only makes our method applicable across different settings but also allows underrepresented modalities to learn more effectively.
3 Methodology
We propose Gradient-Guided Distillation (G2D) to mitigate modality imbalance in multimodal learning through a combination of knowledge distillation and sequential modality prioritization. The proposed G2D adapts to labeled multimodal datasets $\mathcal{D}={(x_{i},y_{i})}_{i=1}^{N}$,where each sample $x_{i}$ consists of $k$-modality inputs, denoted as $x_{i}=(x_{i}^{m_{1}},x_{i}^{m_{2}},\dots,x_{i}^{m_{k}})$, and an associated label $y_{i}$.
[…]
3.1 G2D Loss Function
[…]
(a) CREMA-D
(b) AV-MNIST
(c) VGGSound
(d) UR-Funny
Figure 3: Unimodal teacher confidence scores across multimodal datasets. Each line is the confidence of a specific modality (audio, visual, or text). Modality bias on all datasets, with higher scores for one modality, motivates our use of sequential modality prioritization.
1. Multimodal Student Loss ($\mathcal{L}_{S}$) is a supervised loss to map multimodal inputs $x_{i}$ to the label $y_{i}$:
$\mathcal{L}_{S}=\mathbb{E}_{(x,y)\sim[\mathcal{D}]_{m=1}^{k}}\left[\ell(p,y)\right]$ (1)
where $\ell$ is the cross-entropy loss ($\ell(p,y)\!=\!-\sum_{w=1}^{C}y_{w}log(p)$) for C-class classification tasks with $p\!=\!l_{s}(x;\theta_{s})$ or mean squared error ($\ell(p,y)\!=\!\frac{1}{N}\sum_{i=1}^{N}(p-y)^{2}$) for regression tasks with $p\!=\!\sigma(l_{s}(x;\theta_{s}))$, where $\sigma$ is the sigmoid function.
2. Feature Distillation Loss ($\mathcal{L}_{\text{feat}}$): To prevent the student model from discarding information from weaker modalities, we impose an L2-based feature alignment loss that minimizes the discrepancy between multimodal student and unimodal teachers’ feature representations:
$\mathcal{L}_{\text{feat}}^{m}=\mathbb{E}_{x\sim\mathcal{D}}\left[\|\phi_{s}^{m}(x^{m};\theta_{s}^{m})-\phi_{t}^{m}(x^{m};\theta_{t}^{m})\|^{2}\right]$ (2)
where $\phi_{s}^{m}$ and $\phi_{t}^{m}$ are student and teacher features, respectively, as functions of input $x^{m}$ and encoder parameters.
3. Logits Distillation Loss ($\mathcal{L}_{\text{logit}}$): Logit-based distillation enables the student model to capture class-level relationships and decision boundaries defined by the teacher. We integrate a new logit-based distillation using Kullback-Leibler (KL) divergence [15] that learns the distribution from unimodal teachers to the multimodal student:
$\mathcal{L}_{\text{logit}}^{m}=\mathbb{E}_{x\sim\mathcal{D}}\left[\text{KL}\left(\sigma(l_{t}^{m}(x^{m};\theta_{t}^{m}))\|\sigma(l_{s}(x;\theta_{s}))\right)\right]$ (3)
where $\sigma$ denotes the softmax function.
We define the G2D loss as:
[…]
where $\alpha$ and $\beta$ are weighting coefficients for the feature loss and the logit loss, respectively. This formulation enables the multimodal student model to leverage the strengths of each unimodal teacher model effectively. Feature distillation loss ensures that the student retains modality-specific representations, while logit distillation loss aligns the student’s predictions with teacher distributions, capturing inter-class and intra-class dependencies. Although this new learning objective enables maximization of quality, multimodal models can still be biased to dominant modalities as they optimize all modalities simultaneously. We introduce an adaptive training strategy based on unimodal confidence scores to mitigate modality imbalance.
3.2 Quantifying Modality Confidence
Scoring mechanisms are widely used to measure the contributions of individual modalities in multimodal learning [24, 21]. As unimodal models are outside the bounds of modality imbalance, we utilize their confidence in determining the imbalance ratio. Our scoring module in G2D quantify the confidence $\rho$ of each unimodal teacher $T^{m}$ as the batch-wise average of their softmax function:
$\rho_{t}^{m}=\frac{1}{|\mathcal{B}^{m}|}\sum_{(x_{i}^{m},y_{i}^{m})\in\mathcal{B}^{m}}\text{Softmax}(l_{t}^{m}(x_{i}^{m};\theta^{m}))[y_{i}^{m}],$ (5)
where $|\mathcal{B}^{m}|$ represents the number of $m$ modality data samples in the batch, and $\text{Softmax}(l_{t}^{m}(x_{i}^{m};\theta^{m}))[y_{i}^{m}]$ is the probability assigned to the ground truth label $y_{i}^{m}$ by the teacher model $T^{m}$ for the sample $x_{i}$.
The score $\rho_{t}^{m}$ serves as an indicator of how confident the unimodal teacher $T^{m}$ is in predicting the correct label for the given batch. A higher score indicates greater confidence, signifying that modality $m$ dominates in multimodal training. This information is then used to guide gradient updates in the student model, dynamically adjusting training to mitigate the dominance of any single modality.
3.3 Modulating Gradients with Sequential Modality Prioritization (SMP)
Multimodal datasets often undergo modality overfitting during training [36] and give priority only to dominant modalities, limiting the optimization of weaker modalities. Using the modality scores $\rho^{m}$, we find this trend on all four classification datasets, as shown in Figure 3. It is evident that one modality consistently exhibits higher confidence scores, indicating a modality imbalance and dominance over other modalities in the given downstream task. We also note that this pattern is not grounded to a specific modality. For example, in CREMA-D (Figure 3(a)), the audio teacher has higher scores than the video teacher, and in UR-Funny (Figure 3(d)), the text modality remains dominant over audio and visual inputs. These findings align with previous research that identifies modality bias as a prevalent issue in multimodal datasets [47, 1, 12]. This further leads to modality imbalance in multimodal learning, as analyzed in Figure 1. To address this issue, we hypothesize the following:
Hypothesis 1. Leveraging the confidence scores of unimodal models to determine less confident modalities and sequentially prioritizing them during multimodal training can mitigate modality imbalance.
To test this hypothesis, we propose sequential modality prioritization strategy for the multimodal training in our proposed G2D framework. For each training iteration $q$, we rank confidence scores of all teacher models $\rho_{t}^{m}$. If this ranked list is given as $\pi_{t}$, then $\pi_{t}[1]$ corresponds to the least confident modality, and $\pi_{t}[k]$ corresponds to the most dominant modality. We aim to generate an automatic training schedule process to identify the set of prioritized modalities $\mathcal{M}_{q}$ based on the modality confidence scores $\pi_{t}$, as given in Equation 6.
$\mathcal{M}_{q}=\begin{cases}\{\pi_{t}[1]\}&\text{for }1\leq e\leq\tau_{1}\\ \{\pi_{t}[2]\}&\text{for }\tau_{1}<e\leq\tau_{1}+\tau_{2}\\ \vdots\\ \{\pi_{t}[k-1]\}&\text{for }\sum_{j=1}^{k-2}\tau_{j}<e\leq\sum_{j=1}^{k-1}\tau_{j}\\ \{\pi_{t}[1],\dots,\pi_{t}[k]\}&\text{for }\sum_{j=1}^{k-1}\tau_{j}<e\leq\sum_{j=1}^{k}\tau_{j}\end{cases}$ (6)
where $e$ is the training epoch, and $\tau_{j}$ is the hyperparameter set to denote the number of epochs for optimizing $j$-th prioritized modality, where $j$ is the index of the ranked list $\pi_{t}$. The modulation coefficients $\kappa_{q}^{m}$ for each modality $m$ identifies the modality to optimize in multimodal learning, as given in Equation 7.
$\kappa_{q}^{m}=\begin{cases}1&\text{if modality }m\in\mathcal{M}_{q},\\ 0&\text{otherwise},\end{cases}$ (7)
The gradient update for the parameters $\theta_{q}^{m}$ of modality $m$ for iteration $q$ in the multimodal student model $S$ is then:
[…]
During each epoch range $\tau_{j}$, only the corresponding modality $\pi_{t}(j)$ is assigned $\kappa_{q}^{m}=1$, while all others are set to 00, ensuring that the prioritized modality receives full gradient updates. After the prioritized phases for all less confident modalities, the most confident modality ($\pi_{t}(k)$) is trained jointly with all other modalities, with $\kappa_{q}^{m}=1$ for all $m$. This sequential prioritization strategy allows each modality to lead the learning process in turn, thereby mitigating persistent modality dominance. We summarize the complete multimodal training with G2D in Algorithm 1.
Algorithm 1 Multimodal Learning with G2D
1:Input: $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, unimodal teachers $\{T^{m}\}_{m=1}^{k}$, multimodal student $S$, iterations $q$
2:Initialize student model $S$ with random weights
3:Load pre-trained weights for teacher models $\{T^{m}\}_{m=1}^{k}$
4:for iteration $i=0,\dots,q-1$ do
5:     Sample a fresh mini-batch $(x,y)$ from $\mathcal{D}$
6:     Feed-forward the batch through $S$ to obtain $\{f_{s}^{m}\}_{m=1}^{k}$ and $l_{s}$
7:     Feed-forward each modality of a batch through $\{T^{m}\}_{m=1}^{k}$ to get $\{f_{t}^{m}\}_{m=1}^{k}$ and $\{l_{t}^{m}\}_{m=1}^{k}$
[…]
9:     Calculate $\rho_{t}^{m}$ for each modality using Eq. (5)
10:     Calculate $\kappa^{m}$ using Eq. (7)
[…]
12:end for
13:return Trained multimodal student model $S$
Table 1: Performance comparison on various audio-visual datasets reported in accuracy (%). ”Multi” represents the evaluation of the multimodal student model, while ”Audio” and ”Video” rows indicate the performance of modality-specific encoders within the multimodal model. $T_{a}$ and $T_{v}$ denote unimodal evaluations for the audio and video teacher models, respectively.
Dataset $T^{a}$ $T^{v}$
Joint-
Train
MSES MSLR AGM PMR
OGM-
GE
MLA
MM
Pareto
Recon
Boost
DLMG UMT
G2D
CREMA-D Audio 61.69 - 59.95 54.86 54.86 48.58 49.19 58.60 59.27 65.46 57.71 54.37 61.02 56.45
Video - 76.48 27.42 22.57 26.31 57.85 23.25 49.06 64.91 55.24 65.21 70.89 25.40 72.72
Multi - - 67.47 60.99 64.42 78.48 59.13 72.18 79.70 75.13 79.82 83.62 67.61 85.89
AV-MNIST Audio 42.69 - 16.05 27.50 22.72 38.90 37.60 24.53 42.26 42.11 41.50 41.99 31.55 39.10
Video - 65.44 55.83 63.34 62.92 63.65 58.50 55.85 65.30 65.26 64.28 65.04 64.08 65.09
Multi - - 69.77 70.68 70.62 72.14 71.82 71.08 65.32 72.63 72.47 72.14 72.33 73.03
VGGSound Audio 43.39 - 39.22 39.57 39.10 38.15 26.30 37.96 37.56 42.44 42.35 41.54 42.12 39.43
Video - 32.32 18.70 17.85 18.66 25.65 7.12 22.64 32.02 17.94 18.12 23.65 23.77 29.88
Multi - - 50.97 50.76 50.98 47.11 33.07 51.45 51.65 49.69 50.97 52.74 53.78 53.82
4 Experiments
We evaluate the G2D framework on the basis of the following questions: Q1: How does G2D compare to state-of-the-art methods for addressing modality imbalance and overall multimodal performance in supervised tasks? Q2: How does G2D influence the modality gap in multimodal learning?, Q3: Does G2D enhance feature space alignment of unimodal and multimodal models?, Q4: Which fusion and suppression techniques are the best for G2D?, and Q5: How SMP is influencing the multimodal learning?
4.1 G2D Evaluation
4.1.1 Experimental Setup
Datasets. We choose four multimodal classification datasets and one regression dataset. CREMA-D [3] is an audio-visual dataset for speech emotion recognition with six emotion classes. AV-MNIST [34] is a synthetic dataset with PCA-projected MNIST images and audio spectrograms for ten-digit classes. VGGSound [5] is a large-scale audio-visual dataset with 309 classes of everyday audio events, featuring video clips of 10 seconds each. UR-Funny [13] is a binary classification dataset for humor detection that incorporates text, visual gestures, and acoustic modalities. MIS-ME [28] dataset contains raw soil patch images and corresponding meteorological data for the soil moisture estimation task. To the best of our knowledge, we are the first to evaluate the modality imbalance in a multimodal regression setting and use tabular data for multimodal learning problems.
Baselines. We compare G2D with ten state-of-the-art methods that address modality imbalance, including MSES [9], MSLR [43], AGM [21], PMR [8], OGM-GE [24], MLA [46], MM-Pareto [37], ReconBoost[16], DLMG[42], and UMT [7]. We compare the baseline methods across four datasets that include audio, visual, and text modalities. For the regression task, we compare our approach to MISME [28], which estimates soil moisture from soil patch images and meteorological data. We evaluate the performance of each baseline in both unimodal and multimodal contexts for each modality. Please refer to Sec. A.2 for more detailed baseline descriptions.
Backbone and Hyperparameter Settings. For audio-visual datasets (CREMA-D, VGGSound, and AV-MNIST), we use ResNet-18 [14] as the encoder for both audio and video modalities in the teacher and student models. For the UR-Funny dataset, which involves audio, visual, and text modalities, we use a Transformer-based encoder [33] for each modality. For MIS-ME, we adopt MobileNetV2 [29] as the image feature extractor and use the fully connected neural network proposed in [28] for processing tabular meteorological data. To ensure a fair comparison, we use identical backbone architectures across all baseline models and employ late fusion for training. All models are optimized using SGD with a momentum of 0.9 and a weight decay of 0.0001. We use a batch size of 16 across all experiments and train on a single NVIDIA A10 GPU. More details on experimental settings are provided in Sec. A.4.
4.1.2 Results
[…]
G2D on Two Modalities and Audio-Visual Domain. Table 1 presents the following key observations:
1. Unimodal performances ($T_{a}$ and $T_{v}$) and joint-training reveal that modality imbalance is dataset-dependent. On CREMA-D and VGGSound, video performs well in the unimodal setup but becomes suppressed in multimodal training, while it is the opposite in AV-MNIST as audio is underutilized in multimodal setup. This confirms that modality imbalance is a prevalent issue in multimodal learning, leading to suboptimal fusion in joint-training.
2. DLMG, ReconBoost, and gradient modulation methods (AGM, OGM-GE, MLA, and MMPareto) attempt to reduce modality imbalance and improve fusion. While effective, they do not fully bridge the discrepancy among modalities, as imbalance persists across datasets. G2D surpasses all baselines, demonstrating that SMP ensures balanced optimization and better multimodal integration.
3. To the best of our knowledge, UMT is the only knowledge distillation-based baseline addressing modality imbalance. The results show that the proposed $G^{2}D$ loss that distills knowledge from unimodal teachers with the dynamic training strategy using SMP gives better optimization for weak modalities to outperform UMT across all datasets.
G2D on Three Modalities and Text Domain. Unlike prior approaches [39, 8], G2D is not constrained by the number of modalities. We now analyze results with a three-modality dataset on UR-FUNNY, given in Table 2. In this special case of experiment, we compare our results with baseline models using all combinations of modalities. We find with joint-training that text modality is dominant in all multimodal settings and incorporating all modalities leads to the best multimodal performance. (i) G2D performance: We observe that G2D consistently outperforms methods incorporating adaptive training strategies, such as OGM-GE, MMPareto, and ReconBoost, as well as the KD-based approach UMT, demonstrating its effectiveness in mitigating modality imbalance. (ii) Modality depression: Another interesting finding with more than two modalities (A-V-TXT) is the dominant modality (text) depression across most of our baseline models. We suspect that these models give over-prioritization to weak modalities while not allowing the required optimization for dominant modalities. G2D, on the other hand, treats all modalities fairly with the proposed SMP to reduce modality depression.
Table 2: Accuracy (%) on the UR-Funny dataset across all modality combinations. Unimodal teacher performance for audio, visual, and text are $61.57\%$, $58.25\%$, and $61.77\%$, respectively.
Type
Joint-
Train
OGM-
GE
MM
Pareto
Recon
Boost
UMT
G2D
A-V Audio 57.34 59.76 61.77 60.53 54.63 59.05
Visual 53.92 53.82 55.73 57.87 56.44 58.05
Multi 61.57 61.87 61.27 62.07 60.46 62.98
A-TXT Audio 50.30 54.12 58.15 50.18 55.63 59.86
Text 57.44 58.35 58.45 56.98 57.75 58.85
Multi 62.17 62.47 62.88 61.06 62.47 63.28
V-TXT Visual 49.30 55.33 56.04 55.41 56.34 56.34
Text 51.21 58.95 59.15 50.94 53.82 56.04
Multi 62.07 62.98 61.27 60.07 63.18 63.48
A-V-TXT Audio 55.03 50.30 58.05 51.65 50.70 59.15
Visual 54.93 55.73 56.14 55.26 54.93 55.94
Text 58.25 55.71 58.55 56.25 52.72 58.15
Multi 62.58 63.68 62.88 61.37 63.38 65.49
(a) AV-MNIST: Joint-Train
(b) AV-MNIST: G2D
(c) UR-Funny: Joint-Train
(d) UR-Funny: G2D
Figure 4: Modality gap for AV-MNIST and UR-Funny datasets, with G2D increasing the modality separation compared to joint-training.
G2D on the Multimodal Regression Task. To evaluate the robustness and real-world applicability of G2D, we applied it to a soil moisture estimation task using an in-wild raw soil patch dataset [28], captured via cameras. As given in Table 3, we find that modality imbalance occurs in regression tasks as well. We find that G2D outperforms the baseline method MIS-ME, indicating its effectiveness in multimodal regression tasks.
Table 3: G2D for Soil Moisture Regression Task on MIS-ME Dataset with tabular and image Modalities
Metrics Tabular Image Joint-Train MIS-ME G2D
MAPE 15.49 8.22 14.62 7.52 7.01
R2 0.34 0.76 0.42 0.80 0.82
4.2 Analysis of G2D
Analyzing Modality Gap. Prior work has shown that multimodal learning leads to distinct modality-specific embeddings, with a larger modality gap often correlating with improved performance [23, 32]. Following these insights, we visualize the modality gap in AV-MNIST (audio-visual) and UR-Funny (visual-text) datasets, as shown in Figure 4. Compared to joint-training (Figures 4(a) and 4(c)), G2D (Figures 4(b) and 4(d)) results in a more pronounced modality gap, making modalities more distinguishable in the embedding space. This separation is particularly evident in text-inclusive models, facilitating better feature utilization. These results highlight G2D’s ability to mitigate modality bias and enhance multimodal learning by preserving modality-specific characteristics.
(a) Audio Encoder
(b) Video Encoder
Figure 5: Alignment between unimodal and multimodal features in the audio encoder (Figure 5(a)) and the video encoder (Figure 5(b)).
Analyzing Feature Alignment of G2D. We first analyze the robustness of multimodal features by aligning them with unimodal teacher features for both audio and video encoders in CREMA-D. In this experiment, we use cosine similarity to measure the feature alignment. With Figure 5(a) and Figure 5(b), we show that the alignment of both modalities between unimodal and multimodal features is consistently higher for the G2D compared to a simple joint-training without KD. We consider that better feature alignment in G2D is one of the important factors in improving modality imbalance in multimodal learning.
Figure 6: Analyzing Confidence Ratio of Joint-Training with G2D
Analyzing Modality Imbalance with Confidence Ratio. To quantify the impact of modality imbalance, we compute confidence ratio to measure the weaker modality’s confidence relative to its unimodal teacher. Specifically, we first compute the average confidence score $\rho$ for the weaker modality across the entire dataset in both joint-training and G2D. Then, we normalize these values by dividing them by the corresponding unimodal teacher’s confidence score. A lower confidence ratio indicates that the weaker modality is overshadowed during training, while a higher ratio suggests that the weaker modality in multimodal setting is reaching its performance close to that of its unimodal setup. As shown in Fig. 6, G2D consistently yields higher confidence ratios than joint-training, demonstrating its ability to mitigate modality discrepancy by ensuring a more balanced optimization process and preventing weaker modalities from being suppressed in the multimodal training.
4.3 Ablation Study
Experimental setups of the ablation studies and ablation tables with unimodal performances can be found in Sec. A.5 and Sec. A.6, respectively.
Table 4: Effect of SMP on Different Multimodal Methods
Method SMP CREMA-D AV-MNIST UR-Funny
Joint-Train 67.47 69.77 62.58
80.78 72.51 63.58
UMT 67.61 72.33 63.38
82.39 72.68 64.59
G2D loss 78.63 72.76 63.78
85.89 73.03 65.49
Impact of SMP. The results in Table 4 demonstrate two-pronged observations on SMP with three datasets. (i) SMP integration enhances the performance of both vanilla joint-training and the multimodal KD-based approach UMT. (ii) Incorporating SMP to G2D loss achieves the best overall performance, outperforming all baselines, highlighting its role in mitigating modality imbalance and improving multimodal learning. This confirms the effectiveness and adaptability of SMP, not just in G2D but also in other models.
Table 5: Performance comparison of G2D using different fusion strategies across multiple datasets in terms of accuracy (%).
Fusion CREMA-D AV-MNIST VGGSound UR-Funny
Sum 81.59 72.70 50.67 63.08
Concat 83.60 72.98 53.40 64.49
FiLM [25] 84.27 72.73 48.11 63.48
BiGated [19] 81.32 72.89 46.66 63.38
Cross-Attention [4] 85.35 72.96 53.58 65.09
Late Fusion [11] 85.89 73.03 53.82 65.49
G2D with Various Fusion Modules. Table 5 compares the effect of different fusion strategies on G2D. Out of the traditional fusion methods used for the modality imbalance problem, late fusion consistently achieves the best results. This highlights the effectiveness of leveraging independent unimodal representations while preserving their distinct contributions. However, Cross-Attention based fusion closely follows the performance of late fusion, demonstrating its ability to enhance cross-modal interactions by dynamically attending to relevant features. Concat fusion provides strong performance but falls slightly behind Late Fusion. FiLM and BiGated offer adaptive feature integration yet fail to match the top-performing methods.
Table 6: Partial vs. Complete Modality Suppression in G2D
Type CREMA-D AV-MNIST VGGSound UR-Funny
Partial 81.99 72.83 51.16 63.68
Complete 85.89 73.03 53.82 65.49
Modality Suppression in G2D. We compare two suppression strategies: partial suppression and complete suppression. Partial suppression follows OGM-GE [24], applying gradient modulation with $1-\tanh(x)$, where $x$ is the ratio of modality scores. Complete suppression utilizes SMP, which zeroes out gradients of dominant modalities, allowing the suppressed modality to train to convergence. Table 6 shows that complete suppression consistently outperforms partial suppression across all datasets. This improvement arises from SMP’s ability to shift optimization focus toward the suppressed modality, preventing interference and enabling more effective multimodal feature integration. So, complete suppression significantly mitigates modality imbalance and enhances multimodal learning.
Table 7: Effect of $\alpha$ and $\beta$ in G2D with SMP
Dataset ($\mathbf{\alpha}$, $\mathbf{\beta}$) weights
(0, 0) (0.25, 0.75) (0.5, 0.5) (0.75, 0.25) (1, 0) (0, 1) (1, 1)
CREMA-D 80.78 84.41 84.95 84.81 82.39 84.68 85.89
UR-Funny 63.58 64.79 64.29 64.29 64.59 64.69 65.49
Effect of $\alpha$ and $\beta$ in G2D. $\alpha$ and $\beta$ denote the weighting coefficients of feature loss and logit loss, respectively in the proposed G2D loss. Table 7 evaluates the effect of changing the weightage of feature loss and logit loss on G2D. Assigning full weight to both losses ($\alpha=1,\beta=1$) yields the best overall performance, highlighting their combined importance in multimodal learning. In contrast, removing both losses ($\alpha=0,\beta=0$) significantly reduces performance, confirming their necessity. While different weight combinations impact results, incorporating both losses with higher weight leads to greater improvements across datasets.
5 Conclusion
In this paper, we presented G2D, a simple but effective novel framework designed to tackle modality imbalance in multimodal learning through gradient-guided distillation and sequential modality prioritization. By fusing unimodal and multimodal learning objectives together with knowledge distillation and utilizing confidence scores from unimodal teachers, G2D dynamically prioritizes weaker modalities. With these contributions, G2D ensures each modality contributes effectively during training without being overshadowed by any dominant modalities. Our experimental results, which span multiple classification datasets and a regression task, illustrate that G2D enhances feature alignment, mitigates modality imbalance, and outperforms existing state-of-the-art methods. We believe that the proposed gradient modulation strategy holds great potential to advance balanced learning in complex multimodal scenarios, paving the way for more inclusive and robust AI systems.
Supplementary Material
Appendix A Appendix
A.1 Detailed Dataset Description
A.1.1 Crowd-sourced Emotional Multimodal Actors Dataset (CREMA-D) [3]
CREMA-D is a multimodal dataset designed for emotion recognition research. It contains audio-visual recordings of actors portraying a variety of emotional states, including Anger, Disgust, Fear, Happy, Neutral, and Sad. The dataset features actors from diverse racial and ethnic backgrounds, covering a wide age range, which makes it suitable for studying the interplay between audio and visual emotional expressions. Ratings for emotional intensity and accuracy were gathered from crowd-sourced participants. The dataset is divided into a training set of 6,027 samples, a validation set of 669 samples, and a test set of 745 samples, facilitating robust model training and evaluation.
A.1.2 Audio Visual MNIST (AV-MNIST) [34]
AV-MNIST is a synthetic multimodal dataset designed for audio-visual digit classification. It combines visual MNIST digit images, downsampled using PCA to retain 25% of their original energy, with audio samples of spoken digits from the TIDigits dataset [20]. The audio samples are represented as 112 $\times$ 112 spectrograms and are augmented with noise from the ESC-50 dataset [26]. The dataset consists of 70,000 audio-visual pairs, including 55,000 for training, 10,000 for testing, and 5,000 selected from the training set for validation.
A.1.3 VGGSound [5]
VGGSound is a large-scale audio-visual dataset designed for training and evaluating audio recognition models. It consists of over 200,000 video clips sourced from YouTube, each containing audio-visual correspondence where the sound source is visually present in the video. The dataset includes 310 diverse classes covering various real-world environments, such as people, animals, music, and nature. Each clip is 10 seconds long, ensuring both the audio and visual elements are aligned, making it ideal for audio-visual learning tasks.
A.1.4 UR-Funny [13]
UR-FUNNY is a multimodal dataset created for the task of humor detection, utilizing text, visual gestures, and prosodic acoustic cues. The dataset comprises 1,866 video clips collected from TED Talks featuring diverse speakers and covering 417 different topics. Each clip is labeled with binary humor annotations, with an equal number of humorous and non-humorous samples, ensuring a balanced dataset. The multimodal nature of UR-FUNNY makes it particularly suitable for investigating the relationships among different modalities, offering insights into how text, vision, and audio can jointly contribute to understanding humor in a multimodal learning context.
A.2 Details on Baselines
A.2.1 Modality-Specific Early Stopping (MSES) [9]
MSES is a multimodal learning approach that aims to prevent overfitting by independently managing the learning rate for each modality. Within the MSES framework, each modality’s classifier is treated as a separate task, and early stopping is employed when a modality begins to overfit, while others continue to learn. By formulating a multitask setup, MSES allows for the independent regulation of modality-specific learning progress, ensuring that the stronger modalities do not overshadow the weaker ones during joint training. This method effectively prevents overfitting by identifying and stopping learning for modalities when their validation loss fails to improve, thereby maximizing balanced contributions from each modality and enhancing the overall multimodal learning process.
A.2.2 Modality-Specific Learning Rate (MSLR) [43]
MSLR aims to optimize multimodal late-fusion models by assigning unique learning rates to each modality instead of using a single global learning rate. This approach helps prevent the vanishing gradients issue that can occur when learning rates are not tailored to the specific characteristics of each modality. By assigning modality-specific learning rates, MSLR ensures that each modality contributes effectively to the learning process, ultimately improving the overall performance of the multimodal model.
A.2.3 Adaptive Gradient Modulation (AGM) [21]
AGM addresses modality competition in multimodal models by modulating the participation level of each modality during training. Inspired by Shapley value-based attribution and the OGM-GE algorithm, AGM isolates the contribution of individual modalities and modulates their gradient update intensity accordingly, allowing stronger modalities to be suppressed while amplifying weaker ones. This adaptive strategy applies to all types of fusion architectures, thereby boosting the overall model performance by ensuring a balanced contribution from each modality and mitigating dominance effects that lead to suboptimal joint training outcomes.
A.2.4 Prototypical Modality Rebalance (PMR) [8]
PMR tackles the ”modality imbalance” issue by applying different learning strategies to each modality to ensure more balanced learning. Specifically, PMR uses prototypical cross-entropy (PCE) loss to accelerate the slow-learning modality, allowing it to align more closely with prototypical representations, while also reducing the inhibition from dominant modalities via prototypical entropy regularization (PER). The method effectively exploits features of each modality independently and helps prevent one modality from dominating the learning process, thereby enhancing overall multimodal learning performance.
A.2.5 On-the-fly Gradient Modulation with Generalization Enhancement (OGM-GE) [24]
OGM-GE addresses the issue of under-optimization for specific modalities in multimodal learning by dynamically modulating gradient contributions for each modality. This approach balances the learning pace by modulating gradients of modality-specific coefficients during backpropagation, reducing the dominance of stronger modalities and facilitating better feature exploitation of weaker ones. Additionally, OGM-GE incorporates a generalization enhancement mechanism, adding dynamic Gaussian noise to improve model generalization.
A.2.6 Multimodal Learning with Alternating Unimodal Adaptation (MLA) [46]
MLA addresses the issue of modality dominance by alternating the training focus between modalities rather than using conventional joint optimization. This alternating unimodal adaptation helps avoid interference between modalities, allowing each to reach its full potential while still maintaining cross-modal interactions through a shared head. A gradient modification mechanism is introduced to mitigate ”modality forgetting,” thereby preserving cross-modal information learned during previous iterations. At inference, MLA integrates multimodal information dynamically, using uncertainty-based fusion to manage imbalance across modality-specific contributions effectively.
A.2.7 MMPareto: Boosting Multimodal Learning with Innocent Unimodal Assistance [37]
MMPareto aims to enhance multimodal learning by addressing the gradient conflict that arises between unimodal and multimodal learning objectives. The algorithm uses Pareto integration to align gradient directions across objectives, ensuring a final gradient that benefits all modalities without compromising any. By balancing gradient direction and boosting gradient magnitude, MMPareto improves generalization, providing ”innocent unimodal assistance” to enhance the performance of each modality while maintaining the consistency of multimodal learning.
A.2.8 On Uni-Modal Feature Learning in Supervised Multi-Modal Learning [7]
This paper addresses the problem of insufficient learning of unimodal features in multi-modal learning. The proposed framework consists of two approaches: Uni-Modal Teacher (UMT) and Uni-Modal Ensemble (UME). UMT distills unimodal pre-trained features into a multi-modal model during late-fusion training, ensuring that the representations learned for each modality are preserved effectively while maintaining cross-modal interactions. UME, on the other hand, avoids cross-modal interactions by combining the predictions from unimodal models directly, thus preventing negative interference. To decide which approach to use, they employ an empirical trick explained in the paper. In our experiments, we compare against UMT due to its use of knowledge distillation (KD), which aligns with our proposed approach.
A.2.9 ReconBoost: Boosting Can Achieve Modality Reconcilement [16]
ReconBoost introduces a modality-alternating learning paradigm to mitigate modality competition in multimodal learning. Instead of optimizing all modalities simultaneously, ReconBoost updates each modality separately, ensuring that weaker modalities are not overshadowed by stronger ones. A KL-divergence-based reconcilement regularization is incorporated to maximize diversity between current and past updates, aligning the method with gradient-boosting principles. Unlike traditional boosting, ReconBoost only retains the most recent learner per modality, preventing overfitting and excessive reliance on strong modalities. Additionally, it integrates a memory consolidation regularization to preserve historical modality-specific information and a global rectification scheme to refine joint optimization. Empirical results across multiple benchmarks demonstrate that ReconBoost effectively reconciles modality learning dynamics, leading to improved multimodal fusion performance.
A.2.10 Facilitating Multimodal Classification via Dynamically Learning Modality Gap [42]
Dynamically Learning Modality Gap (DLMG) addresses the modality imbalance problem in multimodal learning by focusing on disparities in category label fitting across different modalities. Unlike prior methods that primarily regulate learning rates or gradient contributions, DLMG leverages contrastive learning to align modality representations and reduce dominance effects. The approach dynamically integrates supervised classification loss and contrastive modality matching loss through either a heuristic strategy or a learning-based optimization strategy that adjusts their relative importance during training. By progressively refining modality alignment while maintaining label supervision, DLMG minimizes performance gaps between dominant and non-dominant modalities, leading to a more balanced and effective multimodal learning process.
A.3 Details on Fusion Techniques
A.3.1 Summation
Summation fusion is a straightforward multimodal integration technique where features from multiple modalities are combined through element-wise addition. Each modality contributes independently, and their respective representations are directly summed without any complex cross-modal interactions. In this approach, the output of each modality-specific encoder is first processed by a fully connected layer to generate unimodal predictions, which are then added together to form a unified representation. This combined output is used to compute a loss, which subsequently updates all components involved, including the encoders and the fully connected layers. Summation fusion’s strength lies in its simplicity and ease of implementation, as it does not require intricate fusion mechanisms. However, it does not explicitly capture inter-modal relationships, potentially limiting its effectiveness in scenarios where richer cross-modal interactions are beneficial.
A.3.2 Concatenation
Concatenation fusion is a common strategy for integrating information from different modalities by concatenating their feature vectors along a specified axis. This method combines feature representations directly, allowing the model to consider information from all modalities together as a single, extended vector. Despite enabling joint perception of multimodal data, it does not explicitly model cross-modal interactions. The concatenated features are passed through a fully connected layer, where the input size equals the sum of all encoder output dimensions, and the output size matches the number of classes. During training, the model uses the resulting fusion output to compute the loss and update all the involved parameters, including those of the individual encoders and the fully connected layer. Concatenation fusion is effective in creating a unified feature representation, but it relies on subsequent layers to extract and learn any interactions between the modalities.
A.3.3 Feature-wise Linear Modulation (FiLM) [25]
FiLM is a sophisticated fusion method that integrates information from multiple modalities by adjusting feature representations in one modality according to the information from another. This modulation approach uses conditional inputs to produce parameters that scale and shift feature activations, enabling the model to dynamically adjust its processing based on context. FiLM works by passing the conditioning modality through a layer that outputs these modulation parameters, which then directly adjust the target modality’s features before they proceed to the next layers in the model. By providing targeted feature modulation, FiLM helps capture cross-modal nuances and allows the model to be more adaptive in multimodal learning tasks that require context-sensitive adjustments.
A.3.4 BiLinear Gated Fusion (BiGated) [19]
BiGated fusion combines bilinear pooling and gating mechanisms to enhance the integration of multiple modalities by capturing their complex interactions. This technique explicitly models cross-modal relationships, providing a more expressive and fine-grained fusion strategy compared to simpler approaches like concatenation or summation. In BiGated fusion, each modality passes through its own fully connected layer, much like summation. However, what sets BiGated apart is its use of a gating mechanism—one modality’s hidden state is processed through an activation function (we use sigmoid) to generate a gated weight, which is then used to modulate the contributions of other modalities. This approach ensures that each modality can dynamically influence how other modalities are represented in the fusion process, allowing for a richer and more adaptive multimodal representation before proceeding to the final classification layers.
A.3.5 Cross-Attention Fusion [4]
Cross-attention fusion enables the dynamic and adaptive integration of multiple modalities by allowing each modality to attend to others through bidirectional or all-directional attention mechanisms. This approach explicitly models inter-modal dependencies, ensuring that each modality can selectively focus on the most relevant features from others. In our implementation, for two-modal cases, modality $X$ attends to modality $Y$ and vice versa, refining their representations based on mutual interactions. For three-modal scenarios, all-directional attention is applied, where each modality interacts not only with one other but also with the third, ensuring comprehensive multimodal integration. The attended representations are normalized to enhance stability and mitigate potential imbalances in feature contributions. Finally, the refined features from all modalities are projected into a unified representation through a fully connected layer. This mechanism effectively captures nuanced cross-modal relationships, allowing the model to leverage complementary modality-specific information for robust multimodal learning.
A.3.6 Late Fusion [11]
Late fusion technique involves independently processing each modality through its respective model or encoder, followed by combining the outputs at a later stage to produce the final prediction. This approach allows each modality to be modeled and optimized in isolation, maintaining the unique properties of each data source. However, it may miss opportunities to exploit early cross-modal interactions that could provide additional benefits during feature learning. In late fusion, each modality-specific encoder is followed by its own fully connected layer, which is trained solely on that modality’s data. The fusion output is computed by averaging the outputs of all unimodal models, ensuring that the fusion occurs only after independent learning is complete. This independence provides flexibility and robustness, especially in scenarios where some modalities may be missing, but limits the ability to deeply integrate multimodal relationships early in the learning process.
A.4 Details on Experimental Setups
A.4.1 Model Architectures
ResNet-18
ResNet-18, a convolutional neural network with 18 layers, belongs to the ResNet family and is renowned for addressing the vanishing gradient problem through residual connections. These residual connections allow information to bypass some layers, which helps stabilize training even in deeper networks. In our experiments, we use ResNet-18 as an encoder for both audio and video modalities across CREMA-D, AV-MNIST, and VGGSound datasets. We used a specific weight initialization strategy: Xavier normal for fully connected layers, Kaiming normal for convolutional layers, and constant initialization for batch normalization layers, which facilitated an effective starting point for network training and ensured stable convergence across multimodal tasks.
Transformer
Transformers are powerful architectures designed for handling sequential data and capturing long-range dependencies through self-attention mechanisms. In our implementation, we employ Transformers as encoders for the audio, video, and text modalities of the UR-Funny dataset. Following the approach described in [21], we utilized a 4-layer Transformer encoder with eight attention heads and a hidden dimension of 768 for each modality in the UR-Funny dataset. Input features were projected to a 768-dimensional embedding using a convolutional layer, ensuring consistency across different modalities. We employed a similar initialization strategy to ResNet-18 to facilitate stable training.
A.4.2 Hyperparameters
We trained our models on 1 Nvidia A10 GPU with a batch size of 16 using the SGD optimizer, with a momentum of 0.9 and a weight decay of $1\times 10^{-4}$. We initialized the learning rate at 0.001 and decayed it by a ratio of 0.1 every 200 epochs. For all experiments, we set the random seed to 999 for reproducibility. We defined the G2D loss function as a weighted sum of student loss, feature loss, and logit loss, where $\alpha$ and $\beta$ are weighting coefficients for the feature loss and logit loss, respectively. We set both $\alpha$ and $\beta$ to 1.0 for all datasets. Additionally, for the logit loss, we used a temperature of 1.0 in the KL Divergence without further softening, effectively utilizing hard logits for the training process.
A.5 Ablation Experimental Setup
We give four ablation studies in the main paper for the proposed G2D: (1) Analyze the impact of SMP in improving multimodal performance; (2) Evaluate G2D across multiple fusion modules; (3) Compare the effectiveness of partial vs complete suppression using SMP; (4) Analyze the impact of feature loss ($\alpha$) and distill loss ($\beta$). We follow the existing approach for conventional methods[24] and deactivate one modality for Late Fusion, FiLM, and BiGated, we deactivate one modality within the multimodal model.
To analyze fusion modules, we employ conventional methods (Sum, Concat, Late Fusion), modulation-based fusion techniques (FiLM, BiGated), and Cross-Attention. To measure unimodal performance in a multimodal setting, we follow [46]. Sum method involved feeding features from the evaluated modality into the corresponding network head prior to summation, and the Concat method used partitioned network heads to measure unimodal performance. For Late Fusion, FiLM, BiGated, and Cross-Attention we deactivate one modality within the multimodal model.
For the modality suppression study, we compared two strategies: partial suppression and complete suppression. Partial suppression was performed using the technique introduced in OGM-GE, which involved gradient modulation with $1-\tanh(x)$, where $x$ is the ratio of scores between two modalities. Complete suppression, on the other hand, applies sequential modality prioritization, allowing each suppressed modality to train to convergence while zeroing out gradients for all other modalities.
A.6 Ablation Tables of Main Paper with Unimodal Performance
Table 8: Performance comparison of G2D with SMP vs. without SMP on three audio-visual datasets reported in accuracy (%).
Type Without SMP With SMP
CREMA-D Audio 51.61 56.45
Visual 56.59 72.72
Multi 78.63 85.89
AV-MNIST Audio 39.62 39.10
Visual 65.18 65.09
Multi 72.76 73.03
VGGSound Audio 42.01 39.43
Visual 28.28 29.88
Multi 52.17 53.82
Table 9: Performance comparison of G2D with SMP vs. without SMP on the UR-Funny dataset across audio-video (A-V), audio-text (A-TXT), visual-text (V-TXT), and audio-visual-text (A-V-TXT) combinations, reported in accuracy (%).
Type Without SMP With SMP
A-V Audio 56.84 59.05
Visual 55.84 58.05
Multi 62.07 62.98
A-TXT Audio 55.73 59.86
Text 59.76 58.85
Multi 63.18 63.28
V-TXT Visual 53.92 56.34
Text 55.53 56.04
Multi 63.38 63.48
A-V-TXT Audio 52.92 59.15
Visual 52.62 55.94
Text 57.14 58.15
Multi 63.78 65.49
Table 10: Performance comparison of G2D using different fusion strategies across multiple datasets in terms of accuracy (%).
Dataset Sum Concat
Late
Fusion
FiLM BiGated
CREMA-D Audio 56.18 51.21 56.45 48.25 48.12
Visual 66.26 71.37 72.72 60.75 63.04
Multi 81.59 83.60 85.89 84.27 81.32
AV-MNIST Audio 39.87 35.73 39.10 41.44 42.40
Video 65.33 64.00 65.09 61.59 65.48
Multi 72.70 72.98 73.03 72.73 72.89
VGGSound Audio 34.96 38.87 39.43 19.09 30.35
Visual 27.28 30.35 29.88 23.12 28.53
Multi 50.67 53.40 53.82 48.11 46.66
UR-Funny Audio 50.70 55.13 59.15 51.41 44.27
Visual 49.30 55.63 55.94 44.27 49.30
Text 58.55 58.15 58.15 52.62 52.21
 Multi 63.08 64.49 65.49 63.48 63.38
Table 11: Partial vs. Complete modality suppression in G2D
Type Partial Complete
CREMA-D Audio 58.60 56.45
Visual 70.30 72.72
Multi 81.99 85.89
AV-MNIST Audio 35.70 39.10
Visual 57.71 65.09
Multi 72.83 73.03
VGGSound Audio 36.54 39.43
Visual 24.76 29.88
Multi 51.16 53.82
UR-Funny Audio 55.03 59.15
Visual 54.93 55.94
Text 58.25 58.15
Multi 63.68 65.49